# Encoded probabilistic imaginary-time evolution on a trapped-ion quantum computer for ground and excited states of spin qubits

### v1

@@ I Introduction
The field of quantum technology has made noteworthy advancements in recent years, with notable developments observed in areas including computing, sensing, and communications. In particular, the application of quantum computers in material calculations offers the promise of performing highly accurate simulations of molecular and material properties, which are crucial for understanding and designing new materials. Hamiltonian simulation [1, 2, 3, 4, 5, 6, 7] enables the modelling of a system’s quantum behaviour, thereby facilitating insights into its dynamics and interactions. Quantum phase estimation (QPE) is indispensable for determining the eigenvalues of Hamiltonians, which are directly related to the ground state energy of the system [8, 9, 10, 11]. Efficient state preparation is a prerequisite for QPE, and a number of approaches have been proposed, including adiabatic time evolution [12, 13, 14, 15], variational quantum eigensolver (VQE) [16, 17], quantum eigenvalue transformation of unitary matrices (QETU) [18], and imaginary-time evolution (ITE) on quantum computers [19, 20, 21, 22, 23, 24, 25]. These algorithms can be used to calculate a variety of physical quantities as key components [26, 27, 28, 29].
@@
Quantinuum H1-1 computer has high-fidelity two-qubit gate of $\sim 2.0\times 10^{-3}$ infidelity with the all-to-all connectivity, performs mid-circuit measurement, and presents quantum volume of $2^{20}$. However, the remaining quantum errors may worsen the computed results, and we address this issue by implementing quantum error detection (QED) [73, 74, 75, 76, 38]. In particular, we employed the Iceberg code [38], which requires only four ancilla qubits for encoding and syndrome measurements. In QED, ancilla qubits are employed for each cycle of syndrome measurement, similarly in PITE, ancilla qubits are utilised for each cycle of the small imaginary time evolution. The feature of the Quantinuum H1-1 computer, which performs mid-circuit measurement and reuse, is desirable for executing QED and PITE.
@@ II.2.1 Probabilistic imaginary-time evolution
In this study, we adopted the PITE method [23] to obtain the lowest energy state among eigenvectors included in an initial state. After determining the ground state, the ground-state energy can be efficiently estimated using QPE [8, 2, 92, 93, 94, 95]. For the Hamiltonian $\mathcal{H}$ for an $n$-qubit system, we define a nonunitary Hermitian operator $\mathcal{M}:=m_{0}e^{-\mathcal{H}\Delta\tau}$, where $\Delta\tau$ denotes a small imaginary time-step size, and an adjustable real parameter $m_{0}$ satisfying $10<m_{0}<$ 1, $m_{0}\neq 1/\sqrt{2}$ is introduced. Introducing an ancilla qubit to embed the nonunitary operator as a submatrix into an extended unitary matrix, we have an entanglement state of the nonunitary operator acted (success state) and other state (failure state) for an input state$|\psi\rangle|0\rangle$, $|\psi\rangle|0\rangle\to\mathcal{M}|\psi\rangle|0\rangle+\sqrt{1-\mathcal{M}^{2}}|\psi\rangle|1\rangle$. Measuring the ancilla qubit as $|0\rangle$ state with probability $\mathbb{P}_{0}=\langle\psi|\mathcal{M}^{2}|\psi\rangle$ collapses the entanglement state into the success state.
@@
A previous study [23] proposed an approximate implementation method for the nonunitary operator $\mathcal{M}$ within the first order of $\Delta\tau$. The approximated ITE operator is given by
@@
where $s_{1}=m_{0}/\sqrt{1-m_{0}^{2}}$ and $s_{1}=\tan\varphi$. The quantum circuit for the approximated PITE is shown in Figure 1. The quantum circuit is comprising of two single-qubit unitary operators $U_{1}=R_{Z}\left(\pi/2\right)R_{Y}(\pi)R_{Z}(-\pi)$ and $U_{2}=R_{Z}\left(\pi/2\right)R_{Y}(\pi)R_{Z}(2\theta_{0}-3\pi/2)$, and forward- and backward-controlled real-time evolution (CRTE) operators, $e^{\Delta\tau s_{1}\mathcal{H}}\otimes|0\rangle\langle 0|+e^{-\Delta\tau s_{1}\mathcal{H}}\otimes|0\rangle\langle 0|$ for time $\Delta\tau s_{1}$, where $\theta_{0}=\kappa\arccos[(m_{0}+\sqrt{1-m_{0}^{2}})/\sqrt{2}]$ and $\kappa=\mathrm{sgn}(m_{0}-1/\sqrt{2})$.
@@ II.2.2 Iceberg code
The Iceberg code encodes $k$ even logical qubits $L:=\{q_{0},q_{1},\ldots,q_{k-1}\}$ into $k+2$ physical qubits $T:=D\cup A$, where $D=L$ and $A:=\{q_{X},q_{Z}\}$. The data register $D$ stores the same quantum state as the logical qubit state of $L$, and the Iceberg code employs two redundant qubits $A$ for encoding. The Iceberg code is classified as stabilizer code, and the code space is stabilized by $S_{X}=\otimes_{i\in T}X_{i}$ and $S_{Z}=\otimes_{i\in T}Z_{i}$. A syndrome measurement using two additional ancilla qubits can detect a single-qubit error. A $k$-bit string state $|x\rangle_{L}=|x_{k-1}\rangle\ldots|x_{1}\rangle|x_{0}\rangle$ is encoded as
@@ II.2.3 Overall circuit
Figure 2: The entire encoded quantum circuit for the PITE comprises eight quantum registers $\{q_{i}\}$ and classical registers $c$. The measurement results with respect to $Z$ basis are stored in the classical registers $c$ (gray lines with arrows). The encoded circuit generates $|0\rangle_{L}$ in code space. The amplitude-encoding circuit is operated to prepare an initial state for the PITE, which contains approximately thirty two-qubit gates. Subsequently, the syndrome measurement is applied. After preparing the encoded initial state, the first step of the PITE is performed, including several syndrome measurements that act at almost equal intervals. Then, $q_{0}$ and $q_{Z}$ registers are observed to collapse the entangled state to the success state. After measuring $q_{0}$ and $q_{Z}$ registers, the re-encode circuit is applied to recover the encoded state, which contains conditional operations based on classical registers. Re-encoding circuits, PITE circuits with syndrome measurement, $Z$-basis measurement are repeatedly performed until sufficient imaginary-time has elapsed.
@@
Although a recent study analyzed that the ITE method on a quantum computer was robust against quantum noise [97], the PITE method usually requires a large number of gate operations. Therefore, quantum devices that do not incorporate quantum error correction (QEC) are undesirable for the PITE method. As a first step, we consider the implementation of the Iceberg code [38]. The entire encoded quantum circuit for the PITE is depicted in Fig. 2. The quantum circuit used in this study consists of a total of eight qubits: six qubits for encoding a four-qubit system and two qubits for syndrome measurement. First, an encoding circuit is applied to generate a logical $|0\rangle_{L}$ state [38]. A single ancilla qubit is used as a flagged qubit to detect errors that occur during encoding [98]. In total, this involves seven two-qubit gate operations. Amplitude encoding is used to generate an encoded initial state $|\psi\rangle_{L}$ for PITE, where approximately 30 physical two-qubit gates are required for logical operations for amplitude encoding. Therefore, we performed syndrome measurements after amplitude encoding. Twelve two-qubit gates are required for the syndrome measurement. The PITE quantum circuit is then executed, as shown in Fig. 3. Because this quantum circuit involves many gate operations, a syndrome measurement is periodically performed within the block. Subsequently, by measuring the $q_{0}$ and $q_{Z}$ quantum registers, the success state is obtained. The logical state is broken by measuring $q_{0}$ and $q_{Z}$ registers. Thus, we reset the $q_{0}$ and $q_{Z}$ registers and perform a re-encoding block to regenerate the encoded state. A detailed circuit diagram of the re-encoding block is presented in Fig. 4, where the flagged qubit is also used. The next PITE block is executed in the re-encoded state. This procedure is repeatedly applied until sufficient imaging time has elapsed. Finally, we performed destructive measurements of the remaining quantum registers.
@@
Figure 3: Th detailed circuit diagram of PITE including syndrome measurements in Figure 2. $U_{1}$ and $U_{2}$ denote the logical single-qubit gates on the ancilla qubit of PITE. Other blocks, excluding the syndrome measurement, are logical RTE operators, and the generators are indicated in these blocks.
@@ III.2.1 Setup
The quantum circuits were constructed using qiskit v1.1.0 [111], where the mapping of the fermionic Hamiltonian to spin operators was performed with qiskit-nature v0.7.2. We then translated it into the quantum circuit of pytket with pytket v1.28.0 [112] and pytket-qiskit v0.53.0, and executed the quantum circuit with pytket-quantinuum v0.33.0. We used the Quantinuum H1-1 trapped-ion computer and its emulator (H1-1E) [68, 38]. The H1-1 quantum computer is composed of 20 qubits and can perform MS gate operations combined with all-to-all connectivity of $\sim 2.0\times 10^{-3}$ infidelity [113]. We ran 1000 shots of quantum computation.
@@ III.2.2 Ground state of the singlet state
Figure 7: Computation results for $\mathrm{(Zr_{Al}V_{N})^{0}}$ complex in w-AlN obtained using a quantum computer. (a) Discard rates of the calculated results owing to errors detected by QED as a function of the number of two-qubit gates. The dashed line denotes the noise model shown in Eq. (25), with $p_{2}=1.6\times 10^{-3}$. (b) Total success probability of PITE according to the PITE steps. (c) Histogram of measured computational basis. Noiseless denotes the noiseless simulation of the PITE, and H1-1E denotes the emulator of the H1-1 quantum computer. All calculations used $N_{\mathrm{shot}}=1000$.
@@
Figure 7(a) shows the results of the discard rate obtained by the QED. The quantum error of H1-1 quantum computer is modeled as depolarizing error and a fitting function of the discard rate is given by [95]
@@
where $N_{\mathrm{2Q}}$ denotes the number of two-quantum gates within the circuit. We adopted the fitting parameter $p_{2}=1.6\times 10^{-3}$ used in a previous experiment [95] and confirmed that it also agrees well with our experiment. It is clearly seen that the discard rates of H1-1 and H1-1E agree well with each other, confirming that the quantum noise that occurs on the actual device can be well modeled by the emulator.
The total success probability for the PITE steps is presented in 7(b). In the PITE with an energy shift, the total success probability approaches the overlap between the initial and ground states. The quantum noise-free calculation result gives a value of 0.78, which is slightly larger than the overlap of 0.75, but this is confirmed to be due to the error in the Trotter-Suzuki decomposition. A reduction in the total success probability was observed in cases without QED.
@@
The measured probabilities of each Slater determinant, that is, the absolute square of the coefficients of the Slater determinant, are shown in Figure 7(c). There is a slight deviation between the results of the FCI calculation and the noiseless simulation of the PITE for the effective Hamiltonian. This deviation is due to the error of the Trotter-Suzuki decomposition and finite number of measurements. In the histogram in Figure 7(c), H1-1 and H1-1E are in good agreement, as in the case of total success probability. In both results, a slight amplitude was observed in states other than $|e_{x\uparrow}e_{x\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle$ and $|e_{y\uparrow}e_{y\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle$. This is considered to be due to the influence of quantum errors that are not detected by the QED, such as two-qubit errors. We quantify the results using classical fidelity [117] defined as
@@
where $|\Psi^{\mathrm{FCI}}\rangle=\sum_{i}c_{i}|i\rangle$ represents the ground state obtained by FCI calculation and $|\Psi\rangle$ is ground state by PITE. The classical fidelity of H1-1 with QED is 0.98, and we can observe that the results have improved compared with the classical fidelity of 0.87, which does not use QED. We note that the unencoded circuit contains 743 two-qubit gates. As a result, this experiment indicates that the influence of quantum errors that are not detected by the QED is small, and the ground state of the singlet state is successfully obtained in a trapped-ion quantum computer.

### v2

@@ I Introduction
The field of quantum technology has made noteworthy advancements in recent years, with notable developments observed in areas including computing, sensing, and communications. In particular, the application of quantum computers in material calculations offers the promise of performing highly accurate simulations of molecular and material properties, which are crucial for understanding and designing new materials. Hamiltonian simulation [1, 2, 3, 4, 5, 6, 7] enables the modelling of a system’s quantum behaviour, thereby facilitating insights into its dynamics and interactions. Quantum phase estimation (QPE) is indispensable for determining the eigenvalues of Hamiltonians, which are directly related to the ground state energy of the system [8, 9, 10, 11]. Efficient state preparation is a prerequisite for QPE, and a number of approaches have been proposed, including adiabatic time evolution [12, 13, 14, 15], variational quantum eigensolver (VQE) [16, 17], quantum eigenvalue transformation of unitary matrices (QETU) [18], and imaginary-time evolution (ITE) on quantum computers [19, 20, 21, 22, 23, 24, 25, 26]. These algorithms can be used to calculate a variety of physical quantities as key components [27, 28, 29, 30].
@@
Quantinuum H1-1 computer has high-fidelity two-qubit gate of $\sim 2.0\times 10^{-3}$ infidelity with the all-to-all connectivity, performs mid-circuit measurement, and presents quantum volume of $2^{20}$. However, the remaining quantum errors may worsen the computed results, and we address this issue by implementing quantum error detection (QED) [74, 75, 76, 77, 39]. In particular, we employed the Iceberg code [39], which requires only four ancilla qubits for encoding and syndrome measurements. While the QED code is unable to reproduce the noiseless results and must therefore discard the erroneous results, it does, however, permit the implementation of the QED code on current computers. In QED, ancilla qubits are employed for each cycle of syndrome measurement, similarly in PITE, ancilla qubits are utilised for each cycle of the small imaginary time evolution. The feature of the Quantinuum H1-1 computer, which performs mid-circuit measurement and reuse, is desirable for executing QED and PITE.
@@ II.2.1 Probabilistic imaginary-time evolution
In this study, we adopted the PITE method [24] to obtain the lowest energy state among eigenvectors included in an initial state. After determining the ground state, the ground-state energy can be efficiently estimated using QPE [8, 2, 93, 94, 95, 96]. For the Hamiltonian $\mathcal{H}$ for an $n$-qubit system, we define a nonunitary Hermitian operator $\mathcal{M}:=m_{0}e^{-\mathcal{H}\Delta\tau}$, where $\Delta\tau$ denotes a small imaginary time-step size, and an adjustable real parameter $m_{0}$ satisfying $10<m_{0}<$ 1, $m_{0}\neq 1/\sqrt{2}$ is introduced. Introducing an ancilla qubit, the nonunitary operator is embeded as a submatrix into an extended unitary matrix as
$\mathcal{U}_{\mathcal{M}}=\begin{pmatrix}\mathcal{M}&\sqrt{1-\mathcal{M}^{2}}\\ \sqrt{1-\mathcal{M}^{2}}&-\mathcal{M}\end{pmatrix},$ (9)
where the unitary matrix $\mathcal{U}_{\mathcal{M}}$ is divided into submatrices depending on the state of the ancilla qubit (the basis of the top left is $|0\rangle\langle 0|$). Operating the extended unitary operator on an input state prepared as $|\psi\rangle|0\rangle$, we have an entanglement state of the nonunitary operator acted (success state) and other state (failure state):
$\mathcal{U}_{\mathcal{M}}|\psi\rangle|0\rangle=\mathcal{M}|\psi\rangle|0\rangle+\sqrt{1-\mathcal{M}^{2}}|\psi\rangle|1\rangle.$ (10)
Measuring the ancilla qubit as $|0\rangle$ state with probability $\mathbb{P}_{0}=\langle\psi|\mathcal{M}^{2}|\psi\rangle$ collapses the entanglement state into the success state, $(1/\sqrt{\mathbb{P}_{0}})\mathcal{M}|\psi\rangle$
@@
Since it is difficult to decompose $\mathcal{U}_{\mathcal{M}}$ for a general Hamiltonian into single and two-qubit gates, a previous study [24] proposed an approximate implementation method for the nonunitary operator $\mathcal{M}$ within the first order of $\Delta\tau$, by executing Taylor expansion. The approximated ITE operator is given by
@@
where $s_{1}=m_{0}/\sqrt{1-m_{0}^{2}}$ and $\varphi=\tan^{-1}s_{1}$. The quantum circuit for the approximated PITE is shown in Figure 1. The quantum circuit is comprising of two single-qubit unitary operators $U_{1}=R_{Z}\left(\pi/2\right)R_{Y}(\pi)R_{Z}(-\pi)$ and $U_{2}=R_{Z}\left(\pi/2\right)R_{Y}(\pi)R_{Z}(2\theta_{0}-3\pi/2)$, and forward- and backward-controlled real-time evolution (CRTE) operators, $e^{\Delta\tau s_{1}\mathcal{H}}\otimes|0\rangle\langle 0|+e^{-\Delta\tau s_{1}\mathcal{H}}\otimes|0\rangle\langle 0|$ for time $\Delta\tau s_{1}$, where $\theta_{0}=\kappa\arccos[(m_{0}+\sqrt{1-m_{0}^{2}})/\sqrt{2}]$ and $\kappa=\mathrm{sgn}(m_{0}-1/\sqrt{2})$.
@@ II.2.2 Iceberg code
In this subsection, we briefly review the implementation of a $\llbracket n+2,n,2\rrbracket$ QED code called the Iceberg code [39]. The Iceberg code encodes $k$ even logical qubits $L:=\{q_{0},q_{1},\ldots,q_{k-1}\}$ into $k+2$ physical qubits $T:=D\cup A$, where $D=L$ and $A:=\{q_{X},q_{Z}\}$. The data register $D$ stores the same quantum state as the logical qubit state of $L$, and the Iceberg code employs two redundant qubits $A$ for encoding. The Iceberg code is classified as stabilizer code, and the code space is stabilized by $S_{X}=\otimes_{i\in T}X_{i}$ and $S_{Z}=\otimes_{i\in T}Z_{i}$. A syndrome measurement using two additional ancilla qubits can detect a single-qubit error. A $k$-bit string state $|x\rangle_{L}=|x_{k-1}\rangle\ldots|x_{1}\rangle|x_{0}\rangle$ is encoded as
@@ II.2.3 Overall circuit
Figure 2: The entire encoded quantum circuit for the PITE comprises eight quantum registers $\{q_{i}\}$ and classical registers $c$. The measurement results with respect to $Z$ basis are stored in the classical registers $c$ (gray lines with arrows). The encoded circuit generates $|0\rangle_{L}$ in code space. The $\{q_{i}\}_{i=0,1,2,3}$ qubits are assigned to the data register, while the remaining qubits are designated as anciila qubits for the iceberg code. The $q_{0}$ qubit is assigned to an ancilla qubit within the PITE circuit. The amplitude-encoding circuit is operated to prepare an initial state for the PITE, which contains approximately thirty two-qubit gates. Subsequently, the syndrome measurement is applied. After preparing the encoded initial state, the first step of the PITE is performed, including several syndrome measurements that act at almost equal intervals. Then, $q_{0}$ and $q_{Z}$ registers are observed to collapse the entangled state to the success state. After measuring $q_{0}$ and $q_{Z}$ registers, the re-encode circuit is applied to recover the encoded state, which contains conditional operations based on classical registers. Re-encoding circuits, PITE circuits with syndrome measurement, $Z$-basis measurement are repeatedly performed until sufficient imaginary-time has elapsed.
@@
Although a recent study analyzed that the ITE method on a quantum computer was robust against quantum noise [98], the PITE method usually requires a large number of gate operations. Therefore, quantum devices that do not incorporate QED and quantum error correction (QEC) are undesirable for the PITE method. As a first step, we consider the implementation of the Iceberg code [39]. The entire encoded quantum circuit for the PITE is depicted in Fig. 2. The quantum circuit used in this study consists of a total of eight qubits: six qubits for encoding a four-qubit system and two qubits ($q_{M0}$ and $q_{M1}$) for syndrome measurement. The $\{q_{i}\}_{i=0,1,2,3}$ qubits are assigned to the data register. Specifically, the $q_{0}$ qubit is assigned to the ancilla qubit within the PITE circuit. First, an encoding circuit is applied to generate a logical $|0\rangle_{L}$ state [39]. A single ancilla qubit is used as a flagged qubit to detect errors that occur during encoding [99]. In total, this involves seven two-qubit gate operations. Amplitude encoding is used to generate an encoded initial state $|\psi\rangle_{L}$ for PITE, where approximately 30 physical two-qubit gates are required for logical operations for amplitude encoding. Therefore, we performed syndrome measurements after amplitude encoding. Twelve two-qubit gates are required for the syndrome measurement. The PITE quantum circuit is then executed, as shown in Fig. 3. Because this quantum circuit involves many gate operations, a syndrome measurement is periodically performed within the block. Subsequently, by measuring the $q_{0}$ and $q_{Z}$ quantum registers, the success state is obtained. The logical state is broken by measuring $q_{0}$ and $q_{Z}$ registers. Thus, we reset the $q_{0}$ and $q_{Z}$ registers and perform a re-encoding block to regenerate the encoded state. A detailed circuit diagram of the re-encoding block is presented in Fig. 4, where the flagged qubit is also used. The next PITE block is executed in the re-encoded state. This procedure is repeatedly applied until sufficient imaging time has elapsed. Finally, we performed destructive measurements of the remaining quantum registers.
@@
Figure 3: The detailed circuit diagram of PITE including syndrome measurements in Figure 2. $U_{1}$ and $U_{2}$ denote the logical single-qubit gates on the ancilla qubit of PITE. Other blocks, excluding the syndrome measurement, are logical RTE operators, and the generators are indicated in these blocks. The logical operation excluding the syndrome measurements, $U_{1}$ and $U_{2}$ in this figure is expressed as $e^{-i\frac{s\Delta\tau}{4}\Lambda}e^{-i\frac{s\Delta\tau}{2}V_{2}}e^{-i\frac{s\Delta\tau}{2}V_{1}}e^{-i\frac{s\Delta\tau}{2}\Lambda_{2}}e^{-i\frac{s\Delta\tau}{2}\Lambda_{1}}e^{-i\frac{s\Delta\tau}{2}V_{2}}e^{-i\frac{s\Delta\tau}{2}V_{1}}e^{-i\frac{s\Delta\tau}{4}\Lambda}$.
@@ III.2.1 Setup
The quantum circuits were constructed using qiskit v1.1.0 [112], where the mapping of the fermionic Hamiltonian to spin operators was performed with qiskit-nature v0.7.2. We then translated it into the quantum circuit of pytket with pytket v1.28.0 [113] and pytket-qiskit v0.53.0, and executed the quantum circuit with pytket-quantinuum v0.33.0. We used the Quantinuum H1-1 trapped-ion computer and its emulator (H1-1E) [69, 39]. The H1-1 quantum computer is composed of 20 qubits and can perform MS gate operations combined with all-to-all connectivity of $\sim 2.0\times 10^{-3}$ infidelity [114]. State preparation and measurement (SPAM) error is about $3.0\times 10^{-3}$. We ran 1000 shots of quantum computation. In Figs. 7 and 8, the results of the actual demonstrations on hardware are represented by H1-1, whereas the other results are based on numerical simulations.
@@
By running our circuit $N_{\mathrm{shot}}$ times on quantum simulator or computer, we evaluate mean estimator $\bar{O}_{\rho}$ by measuring $O$ with respect to quantum state $\rho$ [115]. The deviation of mean estimator $\bar{O}_{\rho}$ from the true value $\mathrm{Tr}[O\rho_{0}]$ obtained from the noiseless state $\rho_{0}$ is assessed by the mean squared error (MSE) given by
$\mathrm{MSE}[\bar{O}_{\rho}]=\mathrm{bias}[\bar{O}_{\rho}]^{2}+\mathrm{var}[\bar{O}_{\rho}],$ (27)
where the bias and the variance is calculated as
$\mathrm{bias}[\bar{O}_{\rho}]=\mathrm{Tr}[O\rho]-\mathrm{Tr}[O\rho_{0}],$
$\mathrm{var}[\bar{O}_{\rho}]=\frac{\mathrm{Tr}[O^{2}\rho]-\mathrm{Tr}[O\rho]^{2}}{N_{\mathrm{cir}}},$ (28)
with the number of available outcomes, $N_{\mathrm{cir}}$. Bias and variance are often called systematic and statistical errors, respectively. In this study, we adopted projection operators as observable $O$.
@@ III.2.2 Ground state of the singlet state
Figure 7: Computation results for $\mathrm{(Zr_{Al}V_{N})^{0}}$ complex in w-AlN obtained using a quantum computer. (a) Discard rates of the calculated results owing to errors detected by QED as a function of the number of two-qubit gates. The dashed line denotes the noise model shown in Eq. (29), with $p_{2}=1.6\times 10^{-3}$. (b) Total success probability of PITE and (c) its MSE according to the PITE steps. (d) Histogram of measured computational basis and (e) its MSE. Noiseless denotes the noiseless simulation of the PITE, and H1-1E denotes the emulator of the H1-1 quantum computer. The results denoted by H1-1 are obtained from actual demonstrations on hardware, whereas the others are based on numerical simulations. The error bars in Figs. (b) and (d) denotes the statistical error. All calculations used $N_{\mathrm{shot}}=1000$.
@@
Figure 7(a) shows the results of the discard rate obtained by the QED. In this study, we employed the QED code to protect against quantum noise, which necessitated discarding a portion of the measurement outcomes. The quantum error of H1-1 quantum computer is modeled as depolarizing error and a fitting function of the discard rate is given by [96]
@@
where $N_{\mathrm{2Q}}$ denotes the number of two-quantum gates within the circuit. We adopted the fitting parameter $p_{2}=1.6\times 10^{-3}$ used in a previous demonstration [96] and confirmed that it also agrees well with our demonstrations. It is clearly seen that the discard rates of H1-1 and H1-1E agree well with each other, confirming that the quantum noise that occurs on the actual device can be well modeled by the emulator. We mention the reasons behind the relatively moderate discard rate of 71%. The PITE circuit contains the several multi-qubit Pauli rotation gates, which are typically decomposed using a combination of several CNOT and single-qubit gates. In the case of four data qubits, the three- and four-qubit rotation gates can be implemented as a single MS gate as given by Eqs. (13), (14), and (15). Therefore, the number of two-qubit gates was not excessive, resulting in a moderate value of the discard rate. We also remark on the applicability of the noise model in Eq. (29) to other hardware platforms. In Refs. [119, 120], the authors investigated the detection correlation matrix using surface code and compared it with a noise model simulation and that obtained using actual hardware composed of transmon qubits. The noise model based on the depolarizing error underpredicts the results, and the inclusion of leakage, crosstalk, and stray interactions improves the results and achieves a result compatible with the experimental result.
The total success probability for the PITE steps is presented in Figs. 7(b). The total success probability of the noiseless result monotonically decreases, approaching the overlap between the initial and ground states. Note that the true value $\mathrm{Tr}[O\rho_{0}]$, which is employed by the Trotter-Suzuki decomposition, is not the same as the exact value that excludes the algorithmic error. In this calculation, the noiseless result gives a value of 0.78, which is slightly larger than the exact value of 0.75. Figure 7(c) shows the MSE of the total success probability, where the MSE of the noiseless simulation is relatively low in most cases. This reason is that the MSE of the noiseless simulation is primarily influenced by statistical error. Considering quantum noise, a reduction in the total success probability from the noiseless one was observed, as presented in the results without QED code [Fig. 7(b)]. The MSE of the results without QED codes monotonically increases, as shown in Fig. 7(c). This is due to the fact that the systematic error based on quantum noise increases in accordance with the PITE step. Note that when $N_{\mathrm{shot}}$ is fixed, the statistical error is reduced in accordance with the reduction in the total success probability from one, as illustrated in Eq. (28) ( the noiseless MSE in Fig. 7(c)). When employing QED code, the total success probabilities become closer to the noiseless result. The statistical error of the results without QED code does not change through the PITE step proceeding, whereas the QED-employed cases increase as the PITE step increases, as shown by the error bars in Fig. 7(b). The reason for the increase in the statistical error is clear from the reduction in $N_{\mathrm{cir}}$ due to the increase in the discard rate, which is calculated as $N_{\mathrm{cir}}=[1-d(N_{\mathrm{2Q}},p_{2})]N_{\mathrm{shot}}$. That is, the QED code has a trade-off of MSE between the reduction of the systematic error and the increase in the statistical error, as given by Eq. (27). In Fig. 7(c), the MSE for H1-1E and H1-1, when employing the QED code, exhibits a relatively low value of MSE in comparison to the MSE of H1-1 without QED code. This indicates that the benefit of reducing system errors is greater than the cost of increasing statistical errors. We found that the iceberg code effectively mitigates the impact of quantum noise on both numerical simulations and actual demonstrations on hardware. To further enhance the precision of the results, an increase in the number of measurements $N_{\mathrm{shot}}$ allows for an improvement in the accuracy of the results in proportion to $O(1/\sqrt{N_{\mathrm{shot}}})$.
@@
The measured probabilities of each Slater determinant, that is, the absolute square of the coefficients of the Slater determinant, are shown in Figure 7(d, e). There is a slight deviation between the results of the FCI calculation and the noiseless simulation of the PITE for the effective Hamiltonian. The deviation between the exact and noiseless values is attributed to two sources: an algorithmic error based on the Trotter-Suzuki decomposition, and a statistical error resulting from the finite number of measurements. In the histogram in Figure 7(d), H1-1 and H1-1E are in good agreement, as in the case of total success probability. In both results, a slight amplitude was observed in states other than $|e_{x\uparrow}e_{x\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle$ and $|e_{y\uparrow}e_{y\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle$. This is considered to be due to the influence of quantum errors that are not detected by the QED, such as two-qubit errors. Furthermore, the MSE for the measured probabilities of each Slater determinant was evaluated, as illustrated in Fig. 7(e). In most cases, the MSE for H1-1 without QED exhibited considerable MSEs. This illustrates the efficacy of the iceberg code. We quantify the results using classical fidelity [121] defined as
@@
where $|\Psi^{\mathrm{FCI}}\rangle=\sum_{i}c_{i}|i\rangle$ represents the ground state obtained by FCI calculation and $|\Psi\rangle$ is ground state by PITE. The classical fidelity of H1-1 with QED is 0.98, and we can observe that the results have improved compared with the classical fidelity of 0.87, which does not use QED. We note that the unencoded circuit contains 743 two-qubit gates. As a result, this demonstration indicates that the influence of quantum errors that are not detected by the QED is small, and the ground state of the singlet state is successfully obtained in a trapped-ion quantum computer. The calculation of physical quantities is a crucial aspect of this study. Although the calculation of energy is useful for assessing the precision of the calculation, it necessitates a deeper circuit when employing QPE [96] or conducting multiple measurements with a change in Pauli basis. In order to calculate the fidelity, it is necessary to apply state tomography. However, both QPE and state tomography must be applied to logical states, which may result in a significant increase in computational cost. In this study, the inability to provide energy was constrained by the limitations of hardware resources. However, the results and discussion, including the physical quantities, will be possible in the future.